# Quark matter, weak interactions, magnetic field, Landau levels, neutrino emission, quark stars, pulsar kicks

### v1

@@ 1 Introduction
Magnetic fields influence neutrino emission through two main mechanisms: (i) by modifying reaction rates and (ii) by introducing anisotropies in the emission pattern. The first directly affects the overall cooling efficiency, while the second may contribute to large-scale dynamical effects such as pulsar kicks Lai and Qian (1998); Arras and Lai (1999); Sagert and Schaffner-Bielich (2008). These effects have been investigated extensively in nuclear matter Baiko and Yakovlev (1999); Riquelme et al. (2005); Potekhin et al. (2015); Potekhin and Chabrier (2018); Dehman et al. (2022); Tambe et al. (2025); Kumamoto and Welch (2025). By contrast, systematic studies of neutrino emission in magnetized quark matter remained comparatively sparse Belyaev and Gvozdev (2017); Ayala et al. (2018, 2024) until very recently Ghosh and Shovkovy (2025a, b). This lack of comprehensive studies constituted a notable gap, particularly in light of the possibility that quark matter may be realized in some of the densest and most strongly magnetized astrophysical environments known.
@@ 1.2 Setup and model assumptions
We consider unpaired two-flavor ($u,d$) quark matter in $\beta$-equilibrium and electrical neutrality at temperatures below the neutrino trapping threshold Prakash et al. (1997), such that $\mu_{\nu}=0$. The chemical potentials satisfy $\mu_{d}=\mu_{u}+\mu_{e}$. Using weak-coupling, low-temperature estimates, one finds $\mu_{u}\simeq 0.8\mu_{d}$ and $\mu_{e}\simeq 0.2\mu_{d}$, corresponding to number densities $n_{u}\simeq 0.5n_{d}$ and $n_{e}\simeq 3\times 10^{-3}n_{d}$. Although the electron density is small (with less than about one electron per $500$ quarks), electrons play a key role in neutrino emission.
We focus on model parameters relevant for quark-matter cores in compact stars. This implies quark chemical potentials $\mu_{f}$ of order $300~\mathrm{MeV}$ ($f=u,d$) and an electron chemical potential $\mu_{e}$ of order $50~\mathrm{MeV}$. We also assume that the temperature is sufficiently low to avoid neutrino trapping, i.e., $T\lesssim 5~\mathrm{MeV}$, ensuring that $T\ll\mu_{e}\ll\mu_{f}$. Generally, the magnetic field is assumed to be strong but realistic, $\sqrt{|eB|}\lesssim 25~\mathrm{MeV}$, corresponding to $B\lesssim 10^{17}~\mathrm{G}$.
@@ 1.3 Kadanoff-Baym transport equation for neutrinos
Without loss of generality, we assume that the quark matter inside the stellar core is spatially uniform on macroscopic scales. Under this assumption, the nonequilibrium neutrino Green functions $G^{\lessgtr}_{\nu}(t,P)$ depend only weakly on time and can be related to the (left-handed) neutrino distribution function as follows Sedrakian and Dieperink (1999); Schmitt et al. (2006); Berdermann et al. (2016):
@@
Here $P=(p_{0},\bm{p})$ denotes the four-momentum of the neutrino, and $p=|\bm{p}|$ is the magnitude of its spatial momentum. Throughout this work, we consider conditions where neutrinos are not trapped in quark matter, so their chemical potential vanishes ($\mu_{\nu}=0$).
@@
Since the macroscopic evolution of dense matter occurs on timescales much longer than those characteristic of weak interactions, the explicit time dependence of the self-energies and Green’s functions on the right-hand side of Eq. (7) may be neglected.
@@ 2 Neutrino emission from direct Urca processes
The neutrino self-energy entering Eq. (8) can be expressed in terms of the electron propagator and the $W$-boson self-energy in momentum space. Although the presence of a background magnetic field modifies the structure of the charged-particle Green functions, the resulting expression for the neutrino self-energy retains a form closely analogous to that of the zero-field case.
To see this, we write the neutrino self-energy in terms of the Fourier-transformed, translationally invariant parts of the electron propagator and the $W$-boson self-energy,
@@
Despite this conventional-looking representation, the background magnetic field is fully encoded in the structure of $\bar{S}_{e}^{\lessgtr}$ and $\bar{\Pi}^{\gtrless}_{\delta\sigma}$.It is important to emphasize, however, that the transverse components of the four-vectors $P$ and $Q$ do not correspond to physical momenta. For charged particles in a magnetic field, only the energy and the momentum component parallel to the field are conserved quantum numbers, while transverse motion is quantized into Landau levels.
The validity of the Fourier representation above follows from the cancellation of gauge-dependent Schwinger phases. In coordinate space, the Green functions of charged particles and the $W$-boson self-energy take the form
where $\Phi(u,u^{\prime})$ is the Schwinger phase and $u=(t,\bm{r})$. Although these phases explicitly break translational invariance at the level of individual Green functions, they cancel exactly in the product, which is the only combination contributing at leading order to the neutrino self-energy. This cancellation restores effective translational invariance and justifies the momentum-space formulation above.
The electron Green functions in momentum space can be written using a spectral representation as follows:
@@
with $E_{e,n}=\sqrt{2n|eB|+p_{z}^{2}+m_{e}^{2}}$, $\ell=1/\sqrt{|eB|}$, and ${\cal P}_{\pm}=(1\pm is_{\perp}\gamma^{1}\gamma^{2})/2$.
@@ 2.1 Neutrino number production rate
where $\xi_{p_{e}}=2p_{e,\perp}^{2}\ell^{2}$. Using the explicit expression for the imaginary part of the $W$-boson self-energy, one finds that the contraction of the lepton and quark tensors, $L_{n,\lambda}^{\delta\sigma}(\bm{p}_{e},\bm{p}_{\nu})\mbox{Im}\left[\Pi^{R}_{\delta\sigma}(Q)\right]$, appearing on the right-hand side of Eq. (17), is proportional to $\left(\bar{P}\cdot P_{\nu}\right)\left(\bar{K}\cdot Y_{e}\right)$. The latter closely resembles the zero-field particle amplitude Iwamoto (1980, 1982). In addition to the four-momenta of the down and up quarks, $\bar{P}$ and $\bar{K}$, and the neutrino four-momentum $P_{\nu}$, this contraction depends on $Y_{e}$, which serves as an effective analog of the electron four-momentum. Unlike in the zero-field case, however, $Y_{e}$ encapsulates all information about the quantized Landau orbits. The explicit components of $Y_{e}$, are given by
@@ 2.4 Numerical results for direct Urca rate
Our main numerical results for the dimensionless ratio $\dot{\mathcal{E}}_{\nu}/\dot{\mathcal{E}}_{\nu}^{(B=0)}$ as a function of the magnetic field are presented in Fig. 3. For illustrative purposes, we use the following values of the chemical potentials: $\mu_{u}=260~\mathrm{MeV}$ and $\mu_{e}=40~\mathrm{MeV}$. It should be noted, however, that the corresponding rates, when normalized to their zero-field values, are expected to depend only weakly on the specific values of the quark chemical potentials. To account for quark Fermi-liquid effects, we take the strong coupling constant to be $\alpha_{s}=0.3$.
In Fig. 3, we show numerical results for three representative temperatures: $T=0.5~\mathrm{MeV}$, $T=1~\mathrm{MeV}$, and $T=2~\mathrm{MeV}$. As expected, the dimensionless ratios at all three temperatures approach unity as $B\to 0$. To cover a broad range of magnetic fields, from approximately $|eB|\simeq 18~\mathrm{MeV}^{2}$ to $|eB|\simeq 5000~\mathrm{MeV}^{2}$, which is equivalent to the field strengths in a range from about $B\simeq 3\times 10^{15}~\mbox{G}$ to $B\simeq 8.5\times 10^{17}~\mbox{G}$, we use a logarithmic scale on the horizontal axis. Note that the squared-energy units of $|eB|$ can be converted to gauss using the relation $B=1.69\times 10^{14}~\mathrm{G}\,|eB|/\mathrm{MeV}^{2}$.
Figure 3: The neutrino energy emission rates for three fixed temperatures: $T=0.5~\mbox{MeV}$, $T=1~\mbox{MeV}$, and $T=2~\mbox{MeV}$. The rates in the LLL approximation are represented by dotted lines. The inset shows a close-up view of the region with the magnetic field strength below $5\times 10^{16}~\mbox{G}$.
@@
Overall, the energy emission rate decreases with increasing magnetic field strength. However, even for fields as strong as $B\simeq 10^{17}~\mbox{G}$, the suppression is limited to only about $20\%$, suggesting that significant observational consequences are unlikely. Interestingly, the emission rate can be substantially enhanced in the lowest Landau level (LLL) regime, although such a regime is rather unlikely to be realized in nature.
Numerical results for the momentum emission rates are shown in Fig. 4 for the same three fixed temperatures: $T=0.5~\mbox{MeV}$, $T=1~\mbox{MeV}$, and $T=2~\mbox{MeV}$. As before, we show the dimensionless ratio $\dot{\mathcal{P}}_{\nu,z}/\dot{\mathcal{E}}_{\nu}^{(B=0)}$ rather than the net longitudinal momentum emission itself. Similar to the energy emission rate, the net momentum exhibits an oscillatory dependence on the magnetic field, with the oscillation peaks becoming increasingly pronounced as the temperature decreases.
Figure 4: The neutrino net momentum emission rates for three fixed temperatures: $T=0.5~\mbox{MeV}$, $T=1~\mbox{MeV}$, and $T=2~\mbox{MeV}$. The rates in the LLL approximation are represented by dotted lines. The inset shows a close-up view of the region with the magnetic field strength below $5\times 10^{16}~\mbox{G}$.
@@
which quantifies the momentum asymmetry of the emitted neutrinos. For representative parameters, $\mu_{e}=40~\mbox{MeV}$ and $T=1~\mbox{MeV}$, this asymmetry reaches only the percent level for magnetic field strengths as large as $B\simeq 3\times 10^{16}~\mbox{G}$. Moreover, even in the extreme-field limit where all occupied electron states reside in the LLL, the asymmetry remains moderate, with a peak value $\eta\lesssim 0.15$. These values are significantly smaller than those predicted by simplified models that estimate neutrino emission asymmetry based on electron spin polarization alone Sagert and Schaffner-Bielich (2008); Ayala et al. (2018, 2024).
@@ 3 Neutrino-antineutrino synchrotron emission
Here we review the derivation of this process using the Kadanoff–Baym formalism, as developed in Ref. Ghosh and Shovkovy (2025b) and briefly summarized in Sec. 1.3. Rather than computing particle amplitudes directly from wave functions Kaminker et al. (1992), the derivation is formulated in terms of the neutrino self-energy. Working entirely with Green functions, this approach provides a conceptually transparent and computationally efficient framework for evaluating neutrino emissivity in dense, magnetized quark matter.

### v2

@@ 1 Introduction
Magnetic fields influence neutrino emission through two main mechanisms: (i) by modifying reaction rates and (ii) by introducing anisotropies in the emission pattern. The first directly affects the overall cooling efficiency, while the second may contribute to large-scale dynamical effects such as pulsar kicks Lai and Qian (1998); Arras and Lai (1999); Sagert and Schaffner-Bielich (2008). These effects have been investigated extensively in nuclear matter Baiko and Yakovlev (1999); Riquelme et al. (2005); Potekhin et al. (2015); Potekhin and Chabrier (2018); Dehman et al. (2022); Tambe et al. (2025); Kumamoto and Welch (2025). By contrast, systematic studies of neutrino emission in magnetized quark matter remained comparatively sparse Xuewen et al. (2005); Belyaev and Gvozdev (2017); Ayala et al. (2018, 2024) until very recently Ghosh and Shovkovy (2025a, b). This lack of comprehensive studies constituted a notable gap, particularly in light of the possibility that quark matter may be realized in some of the densest and most strongly magnetized astrophysical environments known.
@@ 1.2 Setup and model assumptions
We consider unpaired two-flavor ($u,d$) quark matter in $\beta$-equilibrium and electrical neutrality at temperatures below the neutrino trapping threshold Prakash et al. (1997), such that $\mu_{\nu}=0$. The quark and electron chemical potentials satisfy the following relation: $\mu_{d}=\mu_{u}+\mu_{e}$.
To obtain a rough estimate of the relative number densities of quarks and electrons, one may employ the thermodynamic relations for a non-interacting Fermi gas at $B=0$. For a fermionic species $f$ with degeneracy $g_{f}$, mass $m_{f}$, and chemical potential $\mu_{f}$, the number density at temperature $T$ is
$n_{f}=\frac{g_{f}}{2\pi^{2}}\int_{0}^{\infty}\left(\frac{1}{e^{(E_{f}-\mu_{f})/T}+1}-\frac{1}{e^{(E_{f}+\mu_{f})/T}+1}\right)p^{2}\,dp,$ (3)
where $E_{f}=\sqrt{p^{2}+m_{f}^{2}}$. The first term corresponds to particles and the second to antiparticles. For cold dense matter relevant to compact stars ($T\ll\mu_{f}$), the antiparticle contribution is exponentially suppressed and the zero temperature limit provides a good approximation. The corresponding number density relation for a non-interacting Fermi gas reads
$n_{f}=\frac{g_{f}}{6\pi^{2}}\left(\mu_{f}^{2}-m_{f}^{2}\right)^{3/2}\approx\frac{g_{f}\mu_{f}^{3}}{6\pi^{2}},$ (4)
where, in the second expression, the fermion mass has been neglected. This approximation is sufficient for obtaining an order-of-magnitude estimate in the ultrarelativistic regime, $m_{f}\ll\mu_{f}$. For the quarks, the degeneracy factors are $g_{u}=g_{d}=2N_{c}=6$, accounting for two spin states and three color degrees of freedom, whereas for electrons $g_{e}=2$, reflecting spin degeneracy only. Imposing electric charge neutrality,
$\frac{2}{3}n_{u}-\frac{1}{3}n_{d}-n_{e}=0,$ (5)
and solving it together with beta equilibrium relation, $\mu_{d}=\mu_{u}+\mu_{e}$, one finds approximately $\mu_{u}\simeq 0.8\mu_{d}$ and $\mu_{e}\simeq 0.2\mu_{d}$. Since the number density in Eq. (4) scales as $g_{f}\mu_{f}^{3}$, this implies $n_{u}\simeq 0.5\,n_{d}$ and $n_{e}\simeq 3\times 10^{-3}\,n_{d}$ Shovkovy (2005). Although the electron density is small (with less than about one electron per $500$ quarks), electrons play a key role in neutrino emission.
In this review, we adopt representative model parameters appropriate for quark-matter cores in compact stars. Specifically, we consider quark chemical potentials $\mu_{f}$ of order $300~\mathrm{MeV}$ ($f=u,d$) and an electron chemical potential $\mu_{e}$ of order $50~\mathrm{MeV}$. We also assume that the temperature is sufficiently low to avoid neutrino trapping, i.e., $T\lesssim 5~\mathrm{MeV}$, ensuring that $T\ll\mu_{e}\ll\mu_{f}$. Generally, the magnetic field is assumed to be strong but realistic, $\sqrt{|eB|}\lesssim 25~\mathrm{MeV}$, corresponding to $B\lesssim 10^{17}~\mathrm{G}$.
@@
Let us mention in passing that synchrotron $\nu\bar{\nu}$ emission from electrons, unlike the synchrotron emission from quarks or heavy lepton flavors, is mediated by both charged- and neutral-current diagrams, as shown in Fig. 2. As demonstrated in Refs. Yakovlev and Tschaepe (1981); Kaminker et al. (1992), however, these two contributions lead to rate expressions with similar structures. The total result can therefore be expressed as a single contribution, provided that appropriate effective coupling constants $c_{V}^{f}$ and $c_{A}^{f}$ are used, see Table 1.
Figure 2: The charged- and neutral-current diagrams contributing to synchrotron $\nu\bar{\nu}$ emission from electrons.
@@ 1.3 Kadanoff-Baym transport equation for neutrinos
Without loss of generality, we assume that the quark matter inside the stellar core is spatially uniform on macroscopic scales. Under this assumption, the nonequilibrium neutrino Green’s functions $G^{\lessgtr}_{\nu}(t,P)$ depend only weakly on time and can be related to the (left-handed) neutrino distribution function as follows Sedrakian and Dieperink (1999); Schmitt et al. (2006); Berdermann et al. (2016):
@@
Here $P=(p_{0},\bm{p})$ denotes the four-momentum of the neutrino, and $p=|\bm{p}|$ is the magnitude of its spatial momentum. Throughout this work, we consider conditions where neutrinos are not trapped in quark matter, so their chemical potential vanishes ($\mu_{\nu}=0$). To avoid potential confusion, we also note that, throughout this review, the subscript $\nu$ is used exclusively to denote a neutrino and never as a Lorentz index.
@@
Since the macroscopic evolution of dense matter occurs on timescales much longer than those characteristic of weak interactions, the explicit time dependence of the self-energies and Green’s functions on the right-hand side of Eq. (10) may be neglected. In certain astrophysical settings, such as core-collapse supernovae and neutron star mergers, the dynamical timescales may become comparable to those of weak processes Alford et al. (2018); Arras and Weinberg (2019); Most et al. (2024). While this could appear to invalidate our approximation, this is not necessarily the case. If the average temperature and the quark and electron chemical potentials vary sufficiently slowly, it remains justified to neglect the time dependence on the right-hand side of Eq. (10). This assumes, of course, that the other assumptions of our framework, i.e., low temperature and the absence of neutrino trapping, remain valid.
@@ 2 Neutrino emission from direct Urca processes
Although the presence of a background magnetic field modifies the structure of the charged-particle Green’s functions, the resulting expression for the neutrino self-energy in Eq. (11) retains a form closely analogous to that in the zero-field case Sedrakian and Dieperink (1999); Schmitt et al. (2006); Berdermann et al. (2016). This can be understood as follows. In coordinate space, the electron Green’s function and the $W$-boson self-energy can be written as
where $\Phi(u,u^{\prime})$ is the Schwinger phase and $u=(t,\bm{r})$. Formally, the gauge-dependent Schwinger phases break translational invariance, thereby making a conventional momentum representation impossible. However, in the one-loop expression for the neutrino self-energy, which involves the product of the electron Green’s function and the $W$-boson self-energy, the Schwinger phases cancel exactly. The remaining parts are translationally invariant and can therefore be Fourier transformed in the standard way
@@
Substituting these Fourier transforms into the neutrino self-energy expression then leads to Eq. (11).
Thus, despite its conventional appearance similar to the zero-field case, Eq. (11) fully incorporates the effects of the background magnetic field through the Fourier transforms of translationally invariant parts of Green’s functions $\bar{S}_{e}^{\lessgtr}(P)$ and $\bar{\Pi}^{\gtrless}_{\delta\sigma}(Q)$. It should be stressed, however, that the transverse components of the four-momenta $P$ and $Q$ cannot be interpreted as physical momenta. For charged particles in a magnetic field, only the energy and the momentum component parallel to the field are conserved quantum numbers. The transverse dynamics, in contrast, is quantized into discrete Landau levels rather than described by continuous momenta.
The Fourier transform of the electron Green’s functions can be written using a spectral representation as follows:
@@
with $E_{e,n}=\sqrt{2n|eB|+p_{z}^{2}+m_{e}^{2}}$, $\ell=1/\sqrt{|eB|}$, ${\cal P}_{\pm}=(1\pm is_{\perp}\gamma^{1}\gamma^{2})/2$, and $s_{\perp}=\mbox{sign}(eB)$.
@@ 2.1 Neutrino number production rate
where $\xi_{p_{e}}=2p_{e,\perp}^{2}\ell^{2}$. Using the explicit expression for the imaginary part of the $W$-boson self-energy, one finds that the contraction of the lepton and quark tensors, $L_{n,\lambda}^{\delta\sigma}(\bm{p}_{e},\bm{p}_{\nu})\mbox{Im}\left[\Pi^{R}_{\delta\sigma}(Q)\right]$, appearing on the right-hand side of Eq. (20), is proportional to $\left(P\cdot P_{\nu}\right)\left(K\cdot Y_{e}\right)$. The latter closely resembles the zero-field particle amplitude Iwamoto (1980, 1982). In addition to the four-momenta of the down and up quarks, $P$ and $K$, and the neutrino four-momentum $P_{\nu}$, this contraction depends on $Y_{e}$, which serves as an effective analog of the electron four-momentum. Unlike in the zero-field case, however, $Y_{e}$ encapsulates all information about the quantized Landau orbits. The explicit components of $Y_{e}$, are given by
@@ 2.4 Numerical results for direct Urca rate
Our main numerical results for the dimensionless ratio $\dot{\mathcal{E}}_{\nu}/\dot{\mathcal{E}}_{\nu}^{(B=0)}$ as a function of the magnetic field are presented in Fig. 4. For illustrative purposes, we use the following values of the chemical potentials: $\mu_{u}=300~\mathrm{MeV}$ and $\mu_{e}=50~\mathrm{MeV}$. It should be noted, however, that the corresponding rates, when normalized to their zero-field values, are expected to depend only weakly on the specific values of the quark chemical potentials. To account for quark Fermi-liquid effects, we take the strong coupling constant to be $\alpha_{s}=0.3$.
In Fig. 4, we show numerical results for four representative nonzero temperatures: $T=0.25~\mbox{MeV}$, $0.5~\mbox{MeV}$, $1~\mbox{MeV}$, and $2~\mbox{MeV}$, together with the limiting case $T\to 0$. As expected, the dimensionless ratios of the rates approach unity as $B\to 0$. To cover a broad range of magnetic fields, from approximately $|eB|\simeq 50~\mathrm{MeV}^{2}$ to $|eB|\simeq 5000~\mathrm{MeV}^{2}$, which is equivalent to the field strengths in a range from about $B\simeq 8.5\times 10^{15}~\mbox{G}$ to $B\simeq 8.5\times 10^{17}~\mbox{G}$, we use a logarithmic scale on the horizontal axis. Note that the squared-energy units of $|eB|$ can be converted to gauss using the relation $B=1.69\times 10^{14}~\mathrm{G}\,|eB|/\mathrm{MeV}^{2}$.
Figure 4: The neutrino energy emission rates for four nonzero temperatures: $T=0.25~\mbox{MeV}$, $0.5~\mbox{MeV}$, $1~\mbox{MeV}$, and $2~\mbox{MeV}$, together with the limiting case $T\to 0$. The inset shows a close-up view of the region with the magnetic field strength below $5\times 10^{16}~\mbox{G}$.
@@
Overall, the energy emission rate decreases with increasing magnetic field strength. However, even for fields as strong as $B\simeq 10^{17}~\mbox{G}$, the suppression is limited to only about $20\%$, suggesting that significant observational consequences are unlikely. An interesting exception arises in the lowest Landau level (LLL) regime, where the emission rate can be significantly enhanced. Realization of this regime requires extremely strong magnetic fields such that only the lowest Landau level is occupied, i.e., $eB\gtrsim\mu_{e}^{2}$, with $\mu_{e}$ denoting the electron chemical potential. For the representative value $\mu_{e}\simeq 50~\mbox{MeV}$ adopted here, this condition translates into $B\gtrsim 4.2\times 10^{17}~\mbox{G}$. Although such field strengths are not excluded in principle, they lie near the upper end of theoretically plausible magnetic fields in stellar interiors, even when possible amplification mechanisms are taken into account Duncan and Thompson (1992); Broderick et al. (2002).
Numerical results for the momentum emission rates are presented in Fig. 5 for the same four nonzero temperatures, $T=0.25~\mbox{MeV}$, $0.5~\mbox{MeV}$, $1~\mbox{MeV}$, and $2~\mbox{MeV}$, together with the limiting case $T\to 0$. As before, we show the dimensionless ratio $\dot{\mathcal{P}}_{\nu,z}/\dot{\mathcal{E}}_{\nu}^{(B=0)}$ rather than the net longitudinal momentum emission itself. Similar to the energy emission rate, the net momentum exhibits an oscillatory dependence on the magnetic field, with the oscillation peaks becoming increasingly pronounced as the temperature decreases.
Figure 5: The neutrino net momentum emission rates for four nonzero temperatures: $T=0.25~\mbox{MeV}$, $0.5~\mbox{MeV}$, $1~\mbox{MeV}$, and $2~\mbox{MeV}$, together with the limiting case $T\to 0$. The inset shows a close-up view of the region with the magnetic field strength below $5\times 10^{16}~\mbox{G}$.
@@
which quantifies the momentum asymmetry of the emitted neutrinos. For representative parameters, $\mu_{e}=50~\mbox{MeV}$ and $T=1~\mbox{MeV}$, this asymmetry reaches only the percent level for magnetic field strengths as large as $B\simeq 4\times 10^{16}~\mbox{G}$. Moreover, even in the extreme-field limit where all occupied electron states reside in the LLL, the asymmetry remains moderate, with a peak value $\eta\lesssim 0.15$. These values are significantly smaller than those predicted by simplified models that estimate neutrino emission asymmetry based on electron spin polarization alone Sagert and Schaffner-Bielich (2008); Ayala et al. (2018, 2024).
@@ 3 Neutrino-antineutrino synchrotron emission
Here we review the derivation of this process using the Kadanoff–Baym formalism, as developed in Ref. Ghosh and Shovkovy (2025b) and briefly summarized in Sec. 1.3. Rather than computing particle amplitudes directly from wave functions Kaminker et al. (1992), the derivation is formulated in terms of the neutrino self-energy. Working entirely with Green’s functions, this approach provides a conceptually transparent and computationally efficient framework for evaluating neutrino emissivity in dense, magnetized quark matter.